OpenStudy (anonymous):

Steel rods shrink 10% when cooled from furnace temperature to room temperature. If a tie rod is 30in long at furnace temperature how long is the cooled tie rod?

OpenStudy (anonymous):

Once again same thing.... 10%=10/100 you want to find what 10% of 30 is so make that x x/30 cross multiply 30(10)=300 100(x)=100x 300/100=3 x=3

OpenStudy (anonymous):

Now because the tie rod is steel and is going to shrink 10% which you just found out is 3in. simply subtract: 30-3=27 inches